The specific heat capacity of water is 4.18 J/(g °C). In a cup, 100 grams of water was put in the microwave to heat it from a temperature of 25 °C, to right at boiling temperature. What is the enthalpy of this water?
1
Expert's answer
2020-05-19T08:47:29-0400

Solution:

The energy, Q, needed to increase the temperature of an object of mass m, by a temperature ΔT, where the substance's specific heat is C, is given by:

Q = m × C × ΔT


Here,

m = 100 g

C = 4.18 J/(gC)

ΔT = Tfinal - Tinitial = (100 − 25)C = 75C


Inputting:

Q = (100 g) × (4.18 J/(gC) × (75C) = 31350 J

Q = ΔH

ΔH = 31350 J


Answer: The enthalpy of this water is 31350 Joules.